OpenStudy (anonymous):

can any one help me with this X^3-7X-9=0

OpenStudy (immanuelv):

X~3.14092

OpenStudy (anonymous):

u can solve this by trial and improvement i root lies between 3 and 4 because f(3) = 27 -21-9 = -3 and f(4) = 64 -27 -9 = 28 try f(3.1) f(3.2) to get closer to the true root a graph of the function will tell u how many real roots there are - could be only 1 or matbe 3

OpenStudy (anonymous):

of course if u have agraphical calculator u can use that

OpenStudy (anonymous):

ok ....thanks much